OpenStudy (anonymous):

write the equation for the inverse of the function y= arctan (x+ pi/2)

OpenStudy (anonymous):

take tan on both the sides that would be tan y = tan (arc tan x+pi/2) that would simplify by graph tan y = pi-x, now replace x by y and y by x

OpenStudy (anonymous):

thanks so its y=tanx-pi/2 @divu.mkr